OpenStudy (anonymous):

ok, help! 3/4 - (-1/3) = I am drawing another blank I got -5/12.. am i right? anyone?

OpenStudy (anonymous):

you are not right.

OpenStudy (anonymous):

oh no... what did i do wrong?

OpenStudy (anonymous):

x - (-a) = x + a

OpenStudy (anonymous):

so its 13/12?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

YAY! Thanks so much! :)

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

-20/33 * 3/50... know the steps for this one?

OpenStudy (anonymous):

(a / b) * (c / d) = ac / bd

OpenStudy (anonymous):

okie dokie, let me try that. :)

OpenStudy (anonymous):

I got 60/1650... can I simplify this?

OpenStudy (anonymous):

yes. 60 = 10 times 6 = 10 times 3 times 2 1650 = 10 times 165 = 10 times 3 times 55

OpenStudy (anonymous):

-2/55 ???

OpenStudy (anonymous):

I really don't know how to do that that well :(

OpenStudy (anonymous):

yup. -2/55 is right

OpenStudy (anonymous):

YAYA! it just hit me. i got it! thanks to you. I'm almost 30 and havent done this is so long. and im a single momma with 2 kids in the army and never have time to study. this is sooooo much help right now!

OpenStudy (anonymous):

If your not tired of helping yet i have a few more that are stomping me. :)

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

I would help you, but make sure you have really thought about your problem before you post it here. I can give you the answers all day, but it is better if you learn how to apply your math skills to problems, then you wouldnt even need my help. :)

OpenStudy (anonymous):

okie dokie deal! :)

OpenStudy (anonymous):

-13/18 divided by 20/27... dont i just flip? and then multiply?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

a / b = a times 1 / b

OpenStudy (anonymous):

so 13/18 * 27/20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, 13*27 ________ 18*20

OpenStudy (anonymous):

right. remember 18 is 9 times 2 and 27 is 9 times 3

OpenStudy (anonymous):

i thought it was -13/18? you are missing a minus sign

OpenStudy (anonymous):

ok, so that just makes this answer a negative right? im lost now, again..

OpenStudy (anonymous):

yeah. -a = -1 times a.

OpenStudy (anonymous):

so the answer i got is -39/40

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

ok, -14 minus x ------------ 56 x= -14

OpenStudy (anonymous):

huh? how can x = -14?

OpenStudy (anonymous):

i think i got this lets see. -14+14 28 -------- = --- 56 56

OpenStudy (anonymous):

oh hahaha! im sorry x is -14 too.

OpenStudy (anonymous):

so -7 3 --- = -1-- 4 4

OpenStudy (anonymous):

it is - (1 3/4) you have to be careful with that.

OpenStudy (anonymous):

no kidding! that one was very tricky huh?

OpenStudy (anonymous):

no it was not tricky. I meant you have to be careful where you put the minus sign. It should apply to the whole of 1 3/4, so we write -(1 3/4) If it applied to just the 1, then it would be -1 3/4

OpenStudy (anonymous):

cant believe i got it right! your a great teacher!

OpenStudy (anonymous):

glad to be of help :)

OpenStudy (anonymous):

so i would writh it as -(1 3/4) right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

x*y*z*z*z*z* write as an expotetial notation would be: xyz^4

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

-7s^3= -343s is this one right?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

is it\[-7s ^{3}\] or \[(-7s)^{3}\]

OpenStudy (anonymous):

(-7s)^3

OpenStudy (anonymous):

okay, then it is \[-343s ^{3}\]

OpenStudy (anonymous):

remember, you have to cube everything in the brackets, that includes the s. You forgot to do that in your initial answer. You cubed 7 to get 343, but you didnt cube the s.

OpenStudy (anonymous):

ohhhh! ok got cha! And im almost done! yaya! you are helping me sooo much!

OpenStudy (anonymous):

-7x^2 + 14xy +8y^2 +5(-3x^2 + 5xy + 2y^2) OMG!

OpenStudy (anonymous):

okay, what is 3a + 5 a?

OpenStudy (anonymous):

8a

OpenStudy (anonymous):

okay good. so what is 3x^2 -15x^2

OpenStudy (anonymous):

ummm, 9x - 225

OpenStudy (anonymous):

x^2 is just another variable. suppose you put a = x^2, then 3x^2-15x^2 becomes 3a -15a = -12a. but a = x^2 so 3x^2-15x^2 = -12 x^2

OpenStudy (anonymous):

OHHHHH!

OpenStudy (anonymous):

ok let me see what i can come up with.. :)

OpenStudy (anonymous):

still struggling...

OpenStudy (anonymous):

treat all of them like variable. say x^2 = a, xy = b and y^2 = c. write it in terms of a,b,c and then substitute them back as x^2,xy and y^2

OpenStudy (anonymous):

-7a - 3a + 14b + 5b +8c + 2c= like this?

OpenStudy (anonymous):

you forgot to multiply the last three terms by 5.

OpenStudy (anonymous):

-4x^2 + 19xy + 13y^2(5)

OpenStudy (anonymous):

no! no no! -7x^2 + 14xy +8y^2 +5(-3x^2 + 5xy + 2y^2) = -7x^2 + 14xy +8y^2 -15x^2 + 25xy + 10y^2)

OpenStudy (anonymous):

5(a+b+c) = 5a + 5b + 5c

OpenStudy (anonymous):

oh boy.. and now I have to go. :(

OpenStudy (anonymous):

will you be on here at all tomorrow?

OpenStudy (anonymous):

I wish I could stay on all night but these kids have to eat and get baths and to bed. thank you so much! you have been sooooo much help! I am gonna look for ya tomorrow and hope i get lucky!